OpenStudy (anonymous):

i think the answer is 6 if im wrong could u tell me why? or if im right could u tell me so i know! What is the solution to y over 2 – 1 = 2? 2 6 5 3

OpenStudy (johnweldon1993):

That depends...is this \(\huge \frac{y}{2} - 1 = 2\) or \(\huge \frac{y}{2-1} = 2\)

OpenStudy (anonymous):

the first one!

OpenStudy (johnweldon1993):

Then you are absolutely correct

OpenStudy (anonymous):

ok thank u!!